Question 3 Verify using the definition of limit (with ϵ and δ) that lim x→1 (1/2 |x − 1| + 3) = 3, explain all the steps and the overall significance of what you do. [16]

1
Expert's answer
2021-10-20T04:18:28-0400

Let us verify using the definition of limit (with "\varepsilon" and "\delta" ) that "\lim\limits_{ x→1} (\frac{1}{2} |x − 1| + 3) = 3." Let "\varepsilon > 0" be arbitrary. Put "\delta=\varepsilon." Then for each "\varepsilon > 0" there exists "\delta>0" such that if "|x-1|<\delta" then "|(\frac{1}{2} |x − 1| + 3) - 3|=|\frac{1}{2} |x − 1||=\frac{1}{2} |x − 1|<\frac{1}{2}\delta=\frac{1}{2}\varepsilon<\varepsilon." Therefore,

"\lim\limits_{ x→1} (\frac{1}{2} |x − 1| + 3) = 3."